OpenStudy (anonymous):

The position of a car as a function of time is given by the values in table t (seconds) 0 1 2 3 4 5 s (feet) 0 12 16 26 38 56 Find the average velocity for the time period beginning when t = 3 and lasting 2 seconds.

OpenStudy (anonymous):

Can someone walk me through this problem?

OpenStudy (anonymous):

Let's throw out the first three values for t and s because they are irrelevant to the problem. We are left with: t = {3,4,5} s = {26,38,56} Understanding that distance = rate * time, we can solve for rate, which is the average velocity. The car travelled (56 - 26) feet between t = 3 and t = 5 so distance = 30. This took two seconds, so time = 2. Therefore: 30 feet = r * 2 seconds r = 15 feet/second.

OpenStudy (amistre64):

the average velocity is the slope formed by using the data as points on a graph: (3,16) and (5,56)

OpenStudy (amistre64):

to find this "slope" subtract one point from the other: s t (5,56) -(3,16) -------- 2,40 ; and form a fraction by stacking s/t in this case slope = 40/2 = 20

OpenStudy (amistre64):



OpenStudy (anonymous):

Yes but at t = 2, s = 26, not 16. :)

OpenStudy (amistre64):

t=2 is outside of the scope of the question

OpenStudy (anonymous):

whoops, my bad. I meant t = 3.

OpenStudy (amistre64):

ah... i see, the mismatched line messed it up for me :)